06/10/99 Evening shift summary

Angela Bellavance (bellavan@rice.edu)
Fri, 11 Jun 1999 00:06:00 -0500 (CDT)

Shift Summary for eve shift, 10/06/99
(Thursday, June 10)

On Shift: ValP, AngelaB

Also present: Ed, Sydney, JimG, RickK, Elizabeth, SashaG,
PeterSh, RickC

Hours of beam delivered at all intensities: 5
Hours of beam delivered at nominal intensity: 4:30
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1:45

Accounting of lost time: DA debugging, beam studies

========================================================================
Runs ended this shift:

Runs ? to 13505 - debugging runs, not usable data
Run 13506 (1k events) - KTEV_PED
Runs 13507 to 13513 - BEAM_MUON (chamber plateauing)
Run 13514 - BEAM_MUON_4E12 (quick check of nom. HV for DCs)
Runs 13515 to 13520 - beam studies runs, not saved
Run 13521 in progress...

Tapes brought to FCC this shift: 0
(including tapes for online split)

========================================================================
Accesses during shift:

Open access until 18:20
No controlled accesses

========================================================================
Overview of shift:

Debugging from day shift continues until 18:30.
Chambers re-plateaued.
Rick Coleman does target scan and checks primary collimator positions.
Rick Ford begins beam studies.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
? 15:30 Inherit debugging session with no beam.
? 16:46 RickK and Elizabeth finish downstairs. Close-up
checklist done by Angela. Search and Secure
requested.
? 17:00 Kumquat button on DA is coming up red but no error
message is generated. LaDue notified.
Val demands somebody bring him some cake.
? 17:26 4Y kumquat missing FFF word. LaDue working on it.
? 17:34 Verified HV on CA and CsI. All values ok (<10V).
? 17:37 Rebooted pipeline crate 1. We can init. pipeline
now, but not everything is fixed.
? 17:42 Peter does something with the laser timing.(?)
? 18:20 Search and Secure complete. Request muons.
? 18:25 HV for CsI channel 3031 adjusted. Online laser
plot updated - valid for run 13506 on.
? 18:30 Beam arrives!
13506 18:35 Pedestal run started.
18:45 Pedestal run ended.
13507 18:50 Beam muon run to plateau chambers.
to 13513 to 19:45
13514 19:50 Beam muon run at nominal HV.
20:10 Run stopped.
13515 20:15 Beam studies run started, not saved to tape.
RickC doing target scan and collimator check.
~20:25 Beamstops opened.
20:50 Gas checklist done.
20:58 HCC buffer full alarm.
? 21:00 DC1 Y-view HV trips off. Could have been RickC's
study or a beam abort ping. Reset by Angela.
? 21:13 RickK notes the PI+- mass peak is too wide.
Realize need to install new alignment constants.
Done by Sasha. Beam is a bit spotty.
13517 21:28 Start new run b/c of new alignment constants.
Elizabeth modifies regenerator timing. RickC
still working.
22:10 ktev3 runs out of buffer memory. Run stopped.
13518 22:15 New run started. RickC still working.
22:37 On-line plots show weak channel in HA (#23?).
13519 22:51 Start new run to clear histograms. "Clear Histos"
button is apparently not working.
22:55 ktev3 runs out of buffer memory again. Viv
paged, calls in. Ed (& Viv & Sydney) will
continue to look for the source of the problem,
but we will try to continue running for RickC's
benefit.
~23:00 Strange spike on all humidity sensors, settles
back quickly.
13520 23:18 New run started.
23:25 RickC declares he is done. Run ended.
13521 23:52 Run for RickF started.
24:00 Handoff to Jason Hamm.

========================================================================
Follow up Notes on Previous Shift:

- CsI channel 3031 reinstalled and voltage adjusted.

- D. Jensen responded to the page about the regenerator's making noise
when it moved. He said it's always made that noise and everything's ok.

- Sydney has done some run control diagnostic work and sent e-mail to
Taku. Awaiting reply.

- Chambers have been re-plateaued by Sasha.

========================================================================
Items to Follow up on Next Shift:

- ktev3 buffer memory problem will be worked on later (Ed, Viv).

- Still problems to solve in run control.

- SDAQ error for RC6 is not really an error. Ignore for now, Theo has
been notified.

Regards, Angela & Val

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke